1.) Determine k such that the solutions of kx^2 - 4x = 7 are non real complex numbers.
2.) Use the discriminant to determine wheather the solutions of 1,776x^2 - 1,492x + 1,984 = 0 are real numbers.
3.)Find a third degree equation that has a solution set of {2,3, -4}

Answer by stanbon(75887)

1.) Determine k such that the solutions of kx^2 - 4x = 7 are non real complex numbers.
kx^2-4x-7 = 0
a=k, b=-4, c=-7
Discriminant = b^2-4ac = 16-4k*-7 = 16+28k
---------
The solutions are non-real when 16+28k<0:
28k < -16
k < -4/7
===================

2.) Use the discriminant to determine wheather the solutions of 1,776x^2 - 1,492x + 1,984 = 0 are real numbers.
1492^2-4*1776*1984 = -11868272
Not Real because the discriminant is negative
Cheers,
Stan H.
3.)Find a third degree equation that has a solution set of {2,3, -4}